### v1

@@ 1 Introduction
The linearized stability analysis of the endemic equilibrium for the MAXboost model was carried out numerically by [6]. In this paper, we refine the analysis of stability switches on the endemic equilibrium branch when the maximum duration of immunity is varied, and fully characterize the stability changes that occur in the MAXboost model. The main difficulty lies in studying characteristic equations where the delay occurs in multiple ways. Delay systems in which the coefficients in the characteristic equation depend on the time delay $\tau$ only through the exponential term $\exp(-\tau d)$ are easier to study, and the theory in such cases is well developed (see, for example, [18]). One can usually explicitly compute the delay values for which stability switches occur. A systematic approach to studying characteristic equations with delay-dependent parameters has been developed by [4] for one constant delay, and extended to multiple delay systems by [1]. Using this method, we provide geometric and analytical criteria for stability switches with respect to the delay. In addition, this approach provides a rigorous method for finding Hopf bifurcations.
@@
In this paper, we further expand the analysis of the MAXboost model by applying a recently developed graphical user interface that allows to perform the numerical bifurcation analysis of equations with discrete and distributed delays using the package MatCont7p6 for MATLAB [21]. Our analysis captures a richer dynamical behaviour than previous ones, including those from [15], [28] and [10], even in the absence of boosting. In particular, we find a series of catastrophic bifurcations that originate from the switch between supercritical and subcritical Hopf bifurcations when varying the boosting force and the maximal duration of immunity, which can result in drastic changes in the epidemiological situation due to small perturbations of parameters. We find interesting bifurcations on the branches of periodic solutions, including Neimark–Sacker bifurcations and the emergence of stable tori, highlighting how waning and boosting of immunity lead to a very complex dynamical landscape. Most importantly, we identify regions of parameters that exhibit bistability, both of one equilibrum and a periodic solutions and of two distinct periodic solutions for the same parameter values. The existence of coexisting stable periodic solutions with different amplitude and period has important consequences in public health, as perturbations due for instance to the implementation of public health interventions could push the system to a different attractor.
The paper is structured as follows. In Section 2, we briefly summarise the MAXboost model under consideration and formally study the stability switches of the endemic equilibrium depending on the maximal duration of immunity when not boosted. In Section 3, we perform a numerical bifurcation analysis with respect to the length of the immune period and the boosting force. We also focus on the special case of waning immunity but no boosting, showing in this case, too, the existence of multiple stability switches and branches of stable periodic solutions with different amplitude and period. In the last section, we discuss the results and the implications of our findings in the context of public health interventions.
@@ 2 The mathematical model
We consider the model studied by [6] for the spread of an infectious pathogen that confers temporary immunity. Let $S$, $I$ and $R$ be the fraction of susceptible, infected, and immune individuals in the population, respectively, so that $S+I+R=1$. Let $d>0$ be the per capita natural mortality rate of individuals, equal to the total population birth rate. Let $\beta$ and $\gamma$ denote the per capita transmission rate and the recovery rate of infected individuals. Upon recovery, individuals are completely immune to the disease, but they re-enter the susceptible class after time $\tau>0$, unless their immune system is boosted before time $\tau$. If immune individuals enter in contact with the pathogen, their immune system can be boosted to the maximal level at a rate $\beta\nu$, where $\nu\geq 0$ represents the boosting force. Hence, the rate at which individuals re-enter the class $S$ at time $t$ is given by the individuals who had maximal immunity at time $t-\tau$ (either for natural recovery or immune boosting), and their immune system has not been boosted between $t-\tau$ and $t$. This rate is given by
[…]
Since the population is constant, we can write $R=1-S-I$, hence the full model is described by
@@
[…]
@@
We refer to [8] for an equivalent formulation of the model as a partial differential equation, and to [6] for a detailed explanation for how these terms emerge in the model equations.
@@
If $\mathcal{R}_{0}>1$, the disease-free equilibrium is unstable and a unique endemic equilibrium $(S^{*},I^{*})$ exists, with $S^{*}=\mathcal{R}_{0}^{-1}$ and $I^{*}>0$. The endemic equilibrium can be stable or unstable depending on parameters. [6] performed a numerical analysis of the stability boundaries of the endemic equilibrium varying the parameters $\nu$ and $\tau$, with all other parameters fixed so that $\mathcal{R}_{0}>1$, using TRACE-DDE, a MATLAB computational tool [5]. The analysis showed that the equilibrium destabilizes through Hopf bifurcations. However, no information is given about the dynamical behaviour when the endemic equilibrium is unstable.
@@ 2.1 Stability switches of the endemic steady state
In this section, based on the method developed by [4], we study the occurrence of possible stability switches of the endemic equilibrium as a result of increasing the delay.
Recall from [6] that linearizing the system (1) about the endemic equilibrium point $(S^{*},I^{*})$, with $I^{*}\geq 0$, results in the characteristic equation
$\lambda^{3}$ $=-\lambda\beta(\gamma+d)I^{*}-\lambda^{2}(d+\beta I^{*})-\lambda^{2}e^{-\lambda\tau}\nu\mu$ (3)
@@
$W(\lambda,\tau)\equiv P(\lambda,\tau)+Q(\lambda,\tau)e^{-\lambda\tau}=0,$ (5)
@@
$\begin{split}a_{1}(\tau)&=d^{2}+\beta^{2}(I^{*})^{2}-2\beta\gamma I^{*}-\nu^{2}\mu^{2},\\ a_{0}(\tau)&=(\beta I^{*}(\gamma+d))^{2}-2\nu\beta I^{*}(d+\beta I^{*})\mu\sigma-(\sigma-\nu\beta I^{*})^{2}\mu^{2}-2\nu^{2}\mu^{2}\beta I^{*}\sigma.\end{split}$ (8)
@@
$J=\left\{\tau:\tau\geq 0,\ \omega(\tau)\text{ is a positive simple root of }F(\omega,\tau)\right\}.$ (9)
@@
$\tau_{n}(\tau)=\frac{\theta(\tau)+2\pi n}{\omega(\tau)},\quad n\in\mathbb{N}_{0},$ (11)
@@
$S_{n}:J\to\mathbb{R},\quad S_{n}(\tau)=\tau-\tau_{n}(\tau),\quad n\in\mathbb{N}_{0}.$ (12)
@@
$S_{n}(\tau^{*})=0\quad\text{for some }n\in\mathbb{N}_{0},$ (13)
then a simple pair of conjugate pure imaginary roots $\lambda_{\pm}(\tau^{*})=\pm i\omega(\tau^{*})$ of $W(\lambda,\tau)=0$ exists at $\tau=\tau^{*}$ that crosses the imaginary axis. Moreover, the theorem also gives conditions that determine the direction of the crossing, i.e., if $\delta(\tau^{*})>0$ then the pair of eigenvalues crosses the imaginary axis from left to right, and from right to left if $\delta(\tau^{*})<0$, where
$\delta(\tau^{*})=\text{sign }\left\{\frac{d\text{Re}(\lambda)}{d\tau}\mid_{\lambda=i\omega(\tau^{*})}\right\}=\text{sign }\left\{F^{\prime}_{\omega}(\omega(\tau^{*}),\tau^{*})\right\}\text{sign }\left\{\frac{dS_{n}(\tau)}{d\tau}\mid_{\tau=\tau^{*}}\right\}.$ (14)
Let us turn back to our model. We can give explicit conditions on the coefficients of the polynomial $F(\omega,\tau)$ in (7), as functions of $\tau$, such that $F$ admits one or two positive $\omega(\tau)$ roots. In the next proposition, we characterize the set of time delays for which positive roots $\omega^{\pm}(\tau)$ exist.
@@
$J^{-}=\left\{\tau:\tau\geq 0,\ a_{0}(\tau)>0\ \text{and}\ D(\tau)>0\ \text{and}\ a_{1}(\tau)<0\right\}$ (15)
@@
$J^{+}$ $=J^{-}\cup\left\{\tau:\tau\geq 0,\ a_{0}(\tau)<0\right\}\cup\left\{\tau:\tau\geq 0,\ a_{0}(\tau)=0\ \text{and}\ a_{1}(\tau)<0\right\},$ (16)
@@
$S^{*}_{\infty}=\frac{1}{\mathcal{R}_{0}},\quad I^{*}_{\infty}=\frac{d}{\beta}(\mathcal{R}_{0}-1).$ (17)
@@
$F^{\prime}_{\omega}(\omega(\tau),\tau)=2\omega(\tau)(2\omega^{2}(\tau)+a_{1}(\tau)),$ (21)
it is straightforward to see that $F^{\prime}_{\omega}(\omega^{+}(\tau),\tau)>0$ for all $\tau\in J^{+}$ and $F^{\prime}_{\omega}(\omega^{-}(\tau),\tau)<0$ for all $\tau\in J^{-}$. Then we can determine the direction of which eigenvalues cross the imaginary axis, i.e., the sign of $\delta(\tau^{*})$ in (14) only by analyzing the derivative of $S_{n}^{+}(\tau)$ and $S_{n}^{-}(\tau)$ at $\tau=\tau^{*}$, respectively.
@@
The endemic equilibrium may undergo stability changes, and eventually, it becomes stable for any value of the boosting $\nu$.
@@
Case 1. Assume that $J^{-}=\emptyset$, i.e., only one root $\omega^{+}(\tau)$ is feasible for $\tau\in J^{+}=(\tau_{min},\tau_{max})$ (or $J^{+}=[0,\tau_{max}$). We need to show that for all $\tau$ larger than the last zero of $S_{n}^{+}(\tau)$ for some $n\in\mathbb{N}_{0}$, all eigenvalues have passed and remain on the left half of the imaginary axis.
If $\tau_{min}=0$, then from (19) follows that $S_{n}^{+}(\tau_{min})<0$ for all $n\in\mathbb{N}_{0}$. If $\tau_{min}>0$, then it is determined by the conditions $a_{0}(\tau_{min})=0$ and $a_{1}(\tau_{min})>0$, see also Figure 1. Therefore,
$\lim_{\tau\downarrow\tau_{min}}\omega^{+}(\tau)=0^{+}\text{ and }\lim_{\tau\downarrow\tau_{min}}S_{n}^{+}(\tau)=-\infty.$ (22)
@@
$\lim_{\tau\uparrow\tau_{max}}\omega^{+}(\tau)=0^{+}\text{ and }\lim_{\tau\uparrow\tau_{max}}S_{n}^{+}(\tau)=-\infty.$ (23)
This, together with the monotonicity property of $S_{n}^{+}$ in (20), implies that $S_{n}^{+}$ can only have an even number of sign changes for each $n$, so there is an even number of possible stability switches on $J^{+}$. Moreover, since the equilibrium is stable at $\tau=0$, the first switch is toward unstable. From Remark 1, we have that $\delta(\tau^{*})>0\ (<0)$ whenever $(S^{+}_{n})^{\prime}(\tau^{*})>0\ (<0)$, so by the monotonicity property of $S_{n}^{+}$, it is straightforward to see that the last switch is at $\tau=\tau^{*}_{2m}$, the zero of $S_{0}^{+}(\tau)$, and that $\delta(\tau^{*}_{2m})<0$, which completes this part of the proof.
@@
$\lim_{\tau\downarrow\tau_{min}^{-}}\omega^{-}(\tau)=0^{+}\text{ and }\lim_{\tau\downarrow\tau_{min}^{-}}S_{n}^{-}(\tau)=-\infty.$ (24)
@@
$\lim_{\tau\downarrow\tau_{min}}S_{n}^{+}(\tau)=\lim_{\tau\downarrow\tau_{min}}S_{n}^{-}(\tau)\quad\text{for all }n\in\mathbb{N}_{0},$ (26)
@@
In the next section, we will use this technique to compute the stability switches for a given set of pertussis-like parameters and different values of $\nu$.

### v2

@@ 1 Introduction
The linearized stability analysis of the endemic equilibrium for the MAXboost model was carried out numerically by [6]. Here, we refine the analysis of stability switches of the endemic equilibrium when the maximum duration of immunity is varied, and fully characterize the stability changes that occur in the MAXboost model. The main difficulty lies in studying characteristic equations where the delay occurs in multiple ways. Delay systems in which the coefficients in the characteristic equation depend on the time delay $\tau$ only through an exponential term $\exp(-\tau d)$ are easier to study, and the theory in such cases is well developed (see, for example, [18]). One can usually explicitly compute the delay values for which stability switches occur. A systematic approach to studying characteristic equations with delay-dependent parameters has been developed by [4] for one constant delay, and extended to multiple delay systems by [1]. Using this method, we provide geometric and analytical criteria for stability switches with respect to the delay. In addition, this approach provides a rigorous method for finding Hopf bifurcations.
@@
In this paper, we further expand the analysis of the MAXboost model by using a recently developed delay equation importer for MatCont 7p6 for MATLAB [21], that allows to perform the numerical bifurcation analysis of equations with discrete and distributed delays. Our analysis captures a richer dynamical behaviour than previous ones, including those from [15], [29] and [10], even in the absence of boosting. In particular, we find a series of catastrophic bifurcations that originate from the switch between supercritical and subcritical Hopf bifurcations when varying the boosting force and the maximal duration of immunity, which can result in drastic changes in the epidemiological situation due to small perturbations of parameters. We find interesting bifurcations on the branches of periodic solutions, including Neimark–Sacker bifurcations and the emergence of stable tori, highlighting how waning and boosting of immunity lead to a very complex dynamical landscape. Most importantly, we identify regions of parameters that exhibit bistability, both of one equilibrum and a periodic solutions and of two distinct periodic solutions for the same parameter values. The existence of coexisting stable periodic solutions differing in amplitude and period has important consequences in public health, as perturbations due for instance to the implementation of public health interventions could push the system to a different attractor.
The paper is structured as follows. In Section 2, we briefly summarise the MAXboost model under consideration and formally study the stability switches of the endemic equilibrium depending on the maximal duration of immunity when not boosted. In Section 3, we perform a numerical bifurcation analysis with respect to the length of the immune period and the boosting force. We also focus on the special case of waning immunity but no boosting, showing in this case, too, the existence of multiple stability switches and branches of stable periodic solutions with different amplitudes and periods. Finally, we briefly comment on the impact of the expected lifetime on stability. In the last section, we discuss the results and the implications of our findings in the context of public health interventions.
@@ 2 The mathematical model
We consider the model studied by [6] for the spread of an infectious pathogen that confers temporary immunity. Let $S$, $I$ and $R$ be the fraction of susceptible, infected, and immune individuals in the population, respectively, so that $S+I+R=1$. Let $d>0$ be the per capita natural mortality rate of individuals, equal to the total population birth rate. Let $\beta$ and $\gamma$ denote respectively the per capita transmission rate and the recovery rate of infected individuals. Upon recovery, individuals are completely immune to the disease, but they re-enter the susceptible class after time $\tau>0$, unless their immune system is boosted before time $\tau$. If immune individuals enter in contact with the pathogen, their immune system can be boosted to the maximal level at a rate $\beta\nu$, where $\nu\geq 0$ represents the boosting force. Hence, the rate at which individuals re-enter the class $S$ at time $t$ is given by the individuals who had maximal immunity at time $t-\tau$, and their immune system has not been boosted between $t-\tau$ and $t$. Specifically, at time $t-\tau$ there are two cohorts of individuals who acquire maximal immunity: those who recover naturally, at total rate $\gamma I(t-\tau)$, and those whose immunity has been boosted, at rate $\nu\beta I(t-\tau)R(t-\tau)$. The probability that individuals in these cohorts do not die nor receive an immunity boost between $t-\tau$ and $t$ is given by $e^{-d\tau-\nu\beta\int_{-\tau}^{0}I(t+u)\mathop{}\!\mathrm{d}u}$. The total inflow into $S$ at time $t$ is therefore given by
$\left[\gamma I(t-\tau)+\nu\beta I(t-\tau)R(t-\tau)\right]\times\exp\left(-d\tau-\nu\beta\int_{-\tau}^{0}I(t+u)\mathop{}\!\mathrm{d}u\right).$
Since the population is constant, we can write $R=1-S-I$, hence the full model is described by the two equations
@@
$\quad+I(t-\tau)\left(\gamma+\nu\beta\left(1-S(t-\tau)-I(t-\tau)\right)\right)\exp\left(-d\tau-\nu\beta\int_{t-\tau}^{t}I(u)\,\mathop{}\!\mathrm{d}u\right)$
@@
We refer to [8] for an equivalent formulation of the model using partial differential equations.
@@
If $\mathcal{R}_{0}>1$, the disease-free equilibrium is unstable and a unique endemic equilibrium $(S^{*},I^{*})$ exists, with $S^{*}=\mathcal{R}_{0}^{-1}$ and $I^{*}>0$ implicitly defined by the equation
$d(1-S^{*})-\beta I^{*}S^{*}+\left(\gamma+\nu\beta\left(1-S^{*}-I^{*}\right)\right)I^{*}e^{-d\tau-\nu\beta\tau I^{*}}=0.$
The endemic equilibrium can be stable or unstable depending on parameters. [6] performed a numerical analysis of the stability boundaries of the endemic equilibrium varying the parameters $\nu$ and $\tau$, with all other parameters fixed so that $\mathcal{R}_{0}>1$, using TRACE-DDE, a MATLAB computational tool [5]. The analysis shows that the equilibrium destabilizes through Hopf bifurcations. However, no information is given about the dynamical behaviour when the endemic equilibrium is unstable.
@@ 2.1 Stability switches of the endemic steady state
In this section, based on the method developed by [4], we study the occurrence of possible stability switches of the endemic equilibrium as a result of increasing the delay $\tau$.
Recall from [6] that linearizing the system (1) about the endemic equilibrium point $(S^{*},I^{*})$, with $I^{*}=I^{*}(\tau)\geq 0$, results in the characteristic equation
$\lambda^{3}$ $=-\lambda\beta(\gamma+d)I^{*}-\lambda^{2}(d+\beta I^{*})-\lambda^{2}e^{-\lambda\tau}\nu\mu$
@@
$W(\lambda,\tau)\equiv P(\lambda,\tau)+Q(\lambda,\tau)e^{-\lambda\tau}=0,$
@@
$\begin{split}a_{1}(\tau)&=d^{2}+\beta^{2}(I^{*})^{2}-2\beta\gamma I^{*}-\nu^{2}\mu^{2},\\ a_{0}(\tau)&=(\beta I^{*}(\gamma+d))^{2}-2\nu\beta I^{*}(d+\beta I^{*})\mu\sigma-(\sigma-\nu\beta I^{*})^{2}\mu^{2}-2\nu^{2}\mu^{2}\beta I^{*}\sigma.\end{split}$
@@
$J=\left\{\tau:\tau\geq 0,\ \omega(\tau)\text{ is a positive simple root of }F(\omega,\tau)\right\}.$
@@
$\tau_{n}(\tau)=\frac{\theta(\tau)+2\pi n}{\omega(\tau)},\quad n\in\mathbb{N}_{0},$
@@
$S_{n}:J\to\mathbb{R},\quad S_{n}(\tau)=\tau-\tau_{n}(\tau),\quad n\in\mathbb{N}_{0}.$
@@
$S_{n}(\tau^{*})=0\quad\text{for some }n\in\mathbb{N}_{0},$
then a simple pair of conjugate pure imaginary roots $\lambda_{\pm}(\tau^{*})=\pm i\omega(\tau^{*})$ of $W(\lambda,\tau)=0$ exists at $\tau=\tau^{*}$ that crosses the imaginary axis. The theorem also gives conditions that determine the direction of the crossing. In particular, let
$\delta(\tau^{*})=\text{sign }\left\{\frac{\mathop{}\!\mathrm{d}\text{Re}(\lambda)}{\mathop{}\!\mathrm{d}\tau}\mid_{\lambda=i\omega(\tau^{*})}\right\}=\text{sign }\left\{F^{\prime}_{\omega}(\omega(\tau^{*}),\tau^{*})\right\}\text{sign }\left\{\frac{\mathop{}\!\mathrm{d}S_{n}(\tau)}{\mathop{}\!\mathrm{d}\tau}\mid_{\tau=\tau^{*}}\right\}.$ (7)
Then, the pair of eigenvalues crosses the imaginary axis from left to right if $\delta(\tau^{*})>0$, and from right to left if $\delta(\tau^{*})<0$.
Let us return to our model. We can give explicit conditions on the coefficients of the polynomial $F(\omega,\tau)$ in (5), as functions of $\tau$, such that $F$ admits one or two positive $\omega(\tau)$ roots. In the next proposition, we characterize the set of time delays for which positive roots $\omega^{\pm}(\tau)$ exist.
@@
$J^{-}=\left\{\tau:\tau\geq 0,\ a_{0}(\tau)>0\ \text{and}\ D(\tau)>0\ \text{and}\ a_{1}(\tau)<0\right\}$
@@
$J^{+}=J^{-}\cup\left\{\tau:\tau\geq 0,\ a_{0}(\tau)<0\right\}\cup\left\{\tau:\tau\geq 0,\ a_{0}(\tau)=0\ \text{and}\ a_{1}(\tau)<0\right\},$
@@
$S^{*}_{\infty}=\frac{1}{\mathcal{R}_{0}},\quad I^{*}_{\infty}=\frac{d}{\beta}(\mathcal{R}_{0}-1).$
@@
$F^{\prime}_{\omega}(\omega(\tau),\tau)=2\omega(\tau)(2\omega^{2}(\tau)+a_{1}(\tau)),$
it is straightforward to see that $F^{\prime}_{\omega}(\omega^{+}(\tau),\tau)>0$ for all $\tau\in J^{+}$ and $F^{\prime}_{\omega}(\omega^{-}(\tau),\tau)<0$ for all $\tau\in J^{-}$. Then we can determine the direction of which eigenvalues cross the imaginary axis, i.e., the sign of $\delta(\tau^{*})$ in (7), only by analyzing the derivative of $S_{n}^{+}(\tau)$ and $S_{n}^{-}(\tau)$ at $\tau=\tau^{*}$, respectively.
@@
The endemic equilibrium may undergo stability changes, and eventually it becomes stable for any value of the boosting $\nu$.
@@
Case 1. Assume that $J^{-}=\emptyset$, i.e., only one root $\omega^{+}(\tau)$ is feasible for $\tau\in J^{+}=(\tau_{min},\tau_{max})$ (or $J^{+}=[0,\tau_{max})$). We need to show that for all $\tau$ larger than the last zero of $S_{n}^{+}(\tau)$ for some $n\in\mathbb{N}_{0}$, all eigenvalues have passed and remain on the left half of the imaginary axis.
If $\tau_{min}=0$, then from (9) it follows that $S_{n}^{+}(\tau_{min})<0$ for all $n\in\mathbb{N}_{0}$. If $\tau_{min}>0$, then it is determined by the conditions $a_{0}(\tau_{min})=0$ and $a_{1}(\tau_{min})>0$, see also Figure 1. Therefore,
$\lim_{\tau\downarrow\tau_{min}}\omega^{+}(\tau)=0^{+}\text{ and }\lim_{\tau\downarrow\tau_{min}}S_{n}^{+}(\tau)=-\infty.$
@@
$\lim_{\tau\uparrow\tau_{max}}\omega^{+}(\tau)=0^{+}\text{ and }\lim_{\tau\uparrow\tau_{max}}S_{n}^{+}(\tau)=-\infty.$
This, together with the monotonicity property of $S_{n}^{+}$ in (10), implies that $S_{n}^{+}$ can only have an even number of sign changes for each $n$, so there is an even number of possible stability switches on $J^{+}$. Moreover, since the equilibrium is stable at $\tau=0$, the first switch is towards unstable. From Remark 1, we have that $\delta(\tau^{*})>0\ (<0)$ whenever $(S^{+}_{n})^{\prime}(\tau^{*})>0\ (<0)$, so by the monotonicity property of $S_{n}^{+}$, it is straightforward to see that the last switch is at $\tau=\tau^{*}_{2m}$, the zero of $S_{0}^{+}(\tau)$, and that $\delta(\tau^{*}_{2m})<0$, which completes this part of the proof.
@@
$\lim_{\tau\downarrow\tau_{min}^{-}}\omega^{-}(\tau)=0^{+}\text{ and }\lim_{\tau\downarrow\tau_{min}^{-}}S_{n}^{-}(\tau)=-\infty.$
@@
$\lim_{\tau\downarrow\tau_{min}}S_{n}^{+}(\tau)=\lim_{\tau\downarrow\tau_{min}}S_{n}^{-}(\tau)\quad\text{for all }n\in\mathbb{N}_{0},$
@@
In the next section, we will use this technique to compute the stability switches for a given set of parameters inspired by pertussis, and different values of $\nu$.